# On the Convergence of Tâtonnement for Linear Fisher Markets

### v1

@@ 1 Introduction
Market equilibrium (ME) is a central solution concept in economics. It describes a steady state of the market in which supply is equal to demand. In recent year, it has found various applications in resource allocation in large-scale online and physical market places due to its fairness, efficiency in allocating items, and tractability under various real-world-inspired settings. Since ME was first studied by Léon Walras in 1874, extensive work has been undertaken to establish the existence and uniqueness of ME in various market models (Walras, 1874; Arrow and Debreu, 1954; Cole et al., 2017). In this paper, we focus on the Fisher market setting, a well-known market model involving $m$ divisible items which are to be allocated among $n$ buyers. Each item has a fixed supply (usually assumed to be $1$) and each buyer is endowed with a fixed budget of money. In a Fisher market, an equilibrium is a set of prices for the items, and an allocation of items to buyers, such that each buyer is allocated the optimal bundle subject to the prices and her budget, and each item is allocated at its supply.
@@
Alongside the development of market settings and equilibrium solution concepts, market equilibrium computation, by itself or as part of an end-to-end approach to real-world resource allocation problems, has seen a growing interest in various research communities in economics, computer science, and operations research (Scarf et al., 1967; Kantorovich, 1975; Daskalakis et al., 2009; Bateni et al., 2018; Babaioff et al., 2019; Othman et al., 2010; Cole et al., 2017; Birnbaum et al., 2011; Kroer et al., 2019, 2021; Gao and Kroer, 2020; Gao et al., 2021). This is especially so in the past two decades, as the emergence of internet marketplaces, including advertising systems, require approximating or maintaining equilibrium variables in large-scale market settings.
@@
Tâtonnement is a simple, natural, and decentralized price-adjustment process where prices of items are adjusted to reflect buyers’ demands: the price of a good increases if the demand exceeds the supply under current prices, and decreases if the supply exceeds the demand. It can also be viewed as an algorithm for computing market equilibria. Intuitively, if the tâtonnement process is guaranteed to converge to a market equilibrium, then it lends more credence to the equilibrium concept as a solution concept for the economic allocation of goods. Indeed, for various market settings, including Fisher markets with non-linear constant-elasticity-of-substitution (CES) utilities, the convergence of tâtonnement has been established and is well-understood (Cole and Fleischer, 2008; Cheung et al., 2019; Avigdor-Elgrabli et al., 2014). A prominent exception is the simplest, and arguably most prevalent, setting of linear Fisher markets (LFM), where buyers have linear (additive) utilities over items. On the one hand, it has long been known that, for LFMs, tâtonnement does not converge to an equilibrium in the strongest sense of a vanishing metric or norm; see, e.g., Cole and Tao (2019), references therein, and Section 4.4 for examples and further discussions. However, on the other hand, little has been shown regarding the existence and detailed characterizations (e.g. convergence rates) of the approximate convergence (if any) of tâtonnement for this setting.
@@
$\max\,\sum_{i=1}^{n}B_{i}\log\langle v_{i},x_{i}\rangle\ \ {\rm s.t.}\ \sum_{i=1}^{n}x_{i}\leq\mathbf{1},\,x=(x_{i})_{i=1}^{n}\geq 0,$ (EG-Primal)
where $B_{i}\in\mathbb{R},v_{i}\in\mathbb{R}^{m},x_{i}\in\mathbb{R}^{m}$ denotes budget, valuation vector, and allocation vector for buyer $i$, respectively, in an LFM composed of $n$ buyers and $m$ items. It turns out that tâtonnement in LFM is equivalent to the subgradient descent method applied to the dual program of Eq. EG-Primal (See, e.g., Cheung et al. (2013)). This can be shown as follows: by taking the Lagrangian of Eq. EG-Primal, we obtain that the dual problem to EG is to minimize the following function
$\varphi(p)=\max_{x\geq 0}\,\mathcal{L}(x,p)=\max_{x\geq 0}\,\sum_{i=1}^{n}\left(B_{i}\log\langle v_{i},x_{i}\rangle-\sum_{j=1}^{m}p_{j}x_{ij}\right)+\sum_{j=1}^{m}p_{j},$ (1)
which is a max-type nonsmooth convex function w.r.t. $p$ by construction. Every subgradient of $\varphi(p)$ is of the form $\mathbf{1}-\sum_{i}\mathbf{x}^{*}_{i}(p)$, where $x^{*}_{i}(p)$ is a maximizer of $\max_{x_{i}\geq 0}\,(B_{i}\log\langle v_{i},x_{i}\rangle-\sum_{j=1}^{m}p_{j}x_{ij})$. By first-order optimality conditions, for any $p$, $x^{*}_{i}(p)$ is the demand of buyer $i$ under $p$. Thus, tâtonnement (in the form of $p^{t+1}\leftarrow p^{t}+\eta(\sum_{i}x^{*}_{i}(p^{t})-1)$, where $\eta\in\mathbb{R}_{>0}$) is equivalent to applying subgradient descent on $\min_{p\in\mathbb{R}^{m}_{\geq 0}}\varphi(p)$. We will use this relation to establish a linear convergence rate for tâtonnement on LFMs.
@@
We consider a classic form of the tâtonnement process in LFMs, which updates prices as follows (where $[\cdot]_{+}$ denotes thresholding at zero):
$p^{t+1}\leftarrow[p^{t}+\eta z^{t}]_{+},\quad\forall\,t,$ (2)
where $z$ denotes excess demand at time $t$. We give the first positive answers to the question of whether tâtonnement converges in LFM: we show that the prices generated by the process converge at a fast linear rate (i.e. exponentially fast) to a small neighborhood around the true equilibrium prices. Moreover, we show that the size of this neighborhood shrinks to 00 as the stepsize $\eta$ goes to 00. Thus, for any approximation level $\epsilon>0$, we can find an $\eta>0$ such that tâtonnement converges linearly to an $\epsilon$-approximate equilibrium. Our results are achieved by leveraging the connection between tâtonnement and subgradient methods, and from there showing that the dual of the EG convex program satisfies the quadratic growth condition around its optimum. Moreover, we must then show bounded subgradients, which we establish by showing that the tâtonnement process is guaranteed to be bounded above and away from zero for its entire trajectory, assuming a sufficiently small stepsize. As is explained in Section 2, many existing works on tâtonnement use some modified version of the process where prices are generally kept away from zero either through projection to a region bounded away from zero, or by artificially upper-bounding excess demands. Our results show that for LFM, such modifications are not necessary as long as stepsizes are sufficiently small. Finally, we extend our analysis framework to an important variant of LFMs: quasi-linear Fisher markets. For such markets, we again show that tâtonnement achieves a linear convergence rate to approximate equilibrium.
@@ 2 Related Work
Initially, economists studied the convergence of tâtonnement in continuous time. There, Arrow et al. (1959) proved that a continuous-time version of tâtonnement converges to ME for markets satisfying weak gross substitutability (WGS). In the remainder of the literature review, we focus on the discrete-time version of tâtonnement, which is the focus of our paper. Codenotti et al. (2005) considered a discrete version of tâtonnement and proved that it converges to an approximate equilibrium in polynomial time for any exchange economy satisfying WGS. Though their tâtonnement updates are additive and similar to ours, they need to apply tâtonnement on a transformed market and use price projection in order to restrict prices to a region that is bounded and bounded away from zero. Cole and Fleischer (2008) considered a multiplicative tâtonnement with artificially upper-bounded excess demands, and showed polynomial convergence for it on non-linear CES markets satisfying WGS ($10\leq\rho<$ 1). Cheung et al. (2012) extended this analysis to some non-WGS markets.
Cheung et al. (2019) proved that the entropic tâtonnement is equivalent to generalized gradient descent with KL divergence, and thus is guaranteed to reach $\epsilon$ fraction of initial distance to ME measured in a potential function in $\mathcal{O}(\frac{1}{\epsilon})$ iterations for the Leontief Fisher market, and $\mathcal{O}(\log{\frac{1}{\epsilon}})$ iterations for the complementary CES Fisher market ($-\infty<\rho<0$). Avigdor-Elgrabli et al. (2014); Cheung (2014) considered multiplicative tâtonnement and showed convergence to an approximate equilibrium for Fisher markets with nested non-linear CES-Leontief utilities.
@@
Goktas et al. (2021) studied the dual of an Eisenberg-Gale-like convex program from the perspective of the expenditure minimization problem (EMP). Goktas et al. (2023) showed an $\mathcal{O}((1+E^{2})/T^{2})$ convergence for an entropic tâtonnement process on Fisher markets with concave, continuous, nonegative, and $1$-homogeneous  [^1] A utility function $u:\mathbb{R}^{m}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is $\alpha$-homogeneous if $u(ax)=a^{\alpha}u(x)$ for any $a\in\mathbb{R}_{\geq 0}$ and $x\in\mathbb{R}^{m}_{\geq 0}$. (CCNH) utilities, where $E$ denotes the upper bound of the elasticity of demand. Note that in the case of LFM, the elasticity of demand is unbounded, and thus this result gives an infinite upper bound.
@@
Subgradient descent methods (SubGDs) have been extensively studied since their introduction in the 1960s. Historically, SubGD was only known to have a $O(1/\sqrt{T})$ convergence rate for the averaged iterate, where $T$ is the number of iterations. Moreover, this convergence was in terms of the optimality gap, and not in terms of the distance to the optimal solution. For tâtonnement, we are interested in the latter, as we want to know how quickly the prices converge to the market equilibrium prices. Recently, Zamani and Glineur (2023) showed a nearly $O(1/\sqrt{T})$ optimal rate for the last iterate, but again only in terms of optimality gap. It is known, however, that SubGD can achieve a faster convergence rate when the objective satisfies various notions of error-bound conditions. For prior work on SubGDs under error bound conditions, see Johnstone and Moulin (2020) and references therein. Particularly relevant to us, SubGD is known to have linear convergence under the quadratic growth condition (QG) (Nedić and Bertsekas, 2001; Karimi et al., 2016), in terms of the distance to the optimal solution.
@@ 3 Linear Fisher markets and Tâtonnement
We consider a linear Fisher market with $n$ buyers and $m$ divisible items. Each buyer $i$ has a budget $B_{i}>0$ and valuation $v_{ij}\geq 0$ for each item $j$. Let $B=\sum_{i\in[n]}B_{i}$ and $B_{\min}=\min_{i\in[n]}B_{i}$. Denote $v_{i}=(v_{i1},\dots,v_{im})$ for all $i$ and $v$ as the valuation matrix whose $i$-th row is $v_{i}$. An allocation (or bundle) of items $x_{i}\in\mathbb{R}_{\geq 0}^{m}$ gives buyer $i$ a utility of $u_{i}(x_{i})=\langle v_{i},x_{i}\rangle$. Without loss of generality, we assume the following:
@@
$\mathcal{D}_{i}(p):=\operatorname*{\textnormal{Argmax}}_{x_{i}}\left\{\langle v_{i},x_{i}\rangle:\langle p,x_{i}\rangle\leq B_{i},\,x_{i}\geq 0\right\}.$ (Demand)
@@
A Fisher market is said to reach a market equilibrium if allocations $x^{*}_{i}\in\mathbb{R}^{m}_{\geq 0},\;\forall\;i\in[n]$ and prices $p^{*}\in\mathbb{R}^{m}_{\geq 0}$ satisfy
  • Buyer optimality: $x^{*}_{i}\in\mathcal{D}_{i}(p^{*})$ for all $i\in[n]$,
  • Market clearance: $\sum_{i=1}^{n}x^{*}_{ij}=1$ for all $j\in[m]$.
@@
$&\bullet\;\text{Pick}\;x_{i}^{t}\in\mathcal{D}_{i}(p^{t})\quad&\forall\;i\in[n]\\ &\bullet\;z^{t}_{j}=\sum_{i=1}^{n}x_{i}^{t}-1\quad&\forall\;j\in[m]\\ &\bullet\;p^{t+1}_{j}=p^{t}_{j}+\eta^{t}z^{t}_{j}\quad&\forall\;j\in[m]\end{aligned}.$ (3)
Typically, many authors have considered some form of projection to ensure $p^{t+1}\geq 0$, or even such that it is strictly bounded away from zero. We will show that for LFM, it is possible to select $\eta^{t}$ such that the prices are ensured to be strictly bounded away from zero without projection. This is critical for our later convergence results. We will mostly focus on the case where stepsizes are constant, i.e. $\eta^{t}=\eta$ for some $\eta>0$.
As shown in Section 1, Eq. 3 is equivalent to applying subgradient descent on the following non-smooth convex minimization problem:
@@
In particular, starting with some feasible point $p^{0}\in\mathbb{R}^{m}_{\geq 0}$, at time $t$,
$&\bullet\;\text{Pick}\;g(p^{t})\in\partial\varphi(p^{t})\\ &\bullet\;p^{t+1}=\Pi_{\mathbb{R}^{m}_{\geq 0}}\left(p^{t}-\eta g(p^{t})\right)\end{aligned}.$ (5)

### v2

@@ 1 Introduction
Market equilibrium (ME) is a central solution concept in economics. It describes a steady state of the market in which supply is equal to demand. In recent year, it has found various applications in resource allocation in large-scale online and physical market places due to its fairness, efficiency in allocating items, and tractability under various real-world-inspired settings. Since ME was first studied by Léon Walras in 1874, extensive work has been undertaken to establish the existence and uniqueness of ME in various market models (Walras 1874; Arrow and Debreu 1954; Cole et al. 2017). In this paper, we focus on the Fisher market setting, a well-known market model involving $m$ divisible items which are to be allocated among $n$ buyers. Each item has a fixed supply (usually assumed to be $1$) and each buyer is endowed with a fixed budget of money. In a Fisher market, an equilibrium is a set of prices for the items, and an allocation of items to buyers, such that each buyer is allocated the optimal bundle subject to the prices and her budget, and each item is allocated at its supply.
@@
Alongside the development of market settings and equilibrium solution concepts, market equilibrium computation, by itself or as part of an end-to-end approach to real-world resource allocation problems, has seen a growing interest in various research communities in economics, computer science, and operations research (Scarf et al. 1967; Kantorovich 1975; Daskalakis, Goldberg, and Papadimitriou 2009; Bateni et al. 2022; Babaioff, Nisan, and Talgam-Cohen 2019; Othman, Sandholm, and Budish 2010; Cole et al. 2017; Birnbaum, Devanur, and Xiao 2011; Kroer et al. 2019, 2021; Gao and Kroer 2020; Gao, Peysakhovich, and Kroer 2021). This is especially so in the past two decades, as the emergence of internet marketplaces, including advertising systems, require approximating or maintaining equilibrium variables in large-scale market settings.
@@
Tâtonnement is a simple, natural, and decentralized price-adjustment process where prices of items are adjusted to reflect buyers’ demands: the price of a good increases if the demand exceeds the supply under current prices, and decreases if the supply exceeds the demand. It can also be viewed as an algorithm for computing market equilibria. Intuitively, if the tâtonnement process is guaranteed to converge to a market equilibrium, then it lends credence to market equilibrium as a solution concept for the economic allocation of goods. Indeed, for various market settings, including Fisher markets with non-linear constant-elasticity-of-substitution (CES) utilities, the convergence of tâtonnement has been established and is well-understood (Cole and Fleischer 2008; Cheung, Cole, and Devanur 2019; Avigdor-Elgrabli, Rabani, and Yadgar 2014). A prominent exception is the simplest, and arguably most prevalent, setting of linear Fisher markets (LFM), where buyers have linear (additive) utilities over items. It has long been known that, for LFMs, tâtonnement does not converge to an equilibrium in the strictest sense, that is, in terms of a vanishing metric or norm; see, e.g., Cole and Tao (2019) and Section 4 for further discussion and examples. However, whether tâtonnement demonstrates any form of approximate convergence remains largely unexplored.
@@
$\max_{x\in\mathbb{R}^{n\times m}}\,\sum_{i=1}^{n}B_{i}\log\langle v_{i},x_{i}\rangle\ \ {\rm s.t.}\ \sum_{i=1}^{n}x_{i}\leq\mathbf{1},$ (EG-Primal)
where $B_{i}\in\mathbb{R},v_{i}\in\mathbb{R}^{m},x_{i}\in\mathbb{R}^{m}$ denotes budget, valuation vector, and allocation vector for buyer $i$, respectively, in an LFM composed of $n$ buyers and $m$ items. It turns out that tâtonnement in the LFM is equivalent to the subgradient descent method applied to the dual program of Eq. EG-Primal (See, e.g., Cheung, Cole, and Devanur (2013)). This can be shown as follows: by taking the Lagrangian of Eq. EG-Primal, we obtain that the dual problem to EG is to minimize the following function
$\varphi(p)$ $=\max_{x\geq 0}\,\mathcal{L}(x,p)$ (1)
$=\max_{x\geq 0}\,\sum_{i=1}^{n}\left(B_{i}\log\langle v_{i},x_{i}\rangle-\sum_{j=1}^{m}p_{j}x_{ij}\right)+\sum_{j=1}^{m}p_{j},$
which is a max-type nonsmooth convex function w.r.t. $p$ by construction. Every subgradient of $\varphi(p)$ is of the form $\mathbf{1}-\sum_{i}x^{\textup{*}}_{i}(p)$, where $x^{\textup{*}}_{i}(p)$ is a maximizer of $\max_{x_{i}\geq 0}\,(B_{i}\log\langle v_{i},x_{i}\rangle-\sum_{j=1}^{m}p_{j}x_{ij})$. By first-order optimality conditions, for any $p$, $x^{\textup{*}}_{i}(p)$ is the demand of buyer $i$ under $p$. Thus, tâtonnement (in the form of $p^{t+1}\leftarrow p^{t}+\eta(\sum_{i}x^{\textup{*}}_{i}(p^{t})-\mathbf{1})$, where $\eta\in\mathbb{R}_{>0}$) is equivalent to applying subgradient descent on $\min_{p\in\mathbb{R}^{m}_{\geq 0}}\varphi(p)$. We will use this relation to establish a linear convergence rate for tâtonnement in LFMs.
@@
We consider a classic form of the tâtonnement process in LFMs, which updates prices as follows:
$p^{t+1}\leftarrow p^{t}+\eta z^{t},\quad\forall\,t,$ (2)
where $z$ denotes excess demand at time $t$. We give the first positive answers to the question of whether tâtonnement converges in LFMs: we show that the prices generated by the process converge at a fast linear rate (i.e. exponentially fast) to a small neighborhood around the true equilibrium prices. Moreover, we show that the size of this neighborhood shrinks to 00 as the stepsize $\eta$ goes to 00. Thus, for any approximation level $\epsilon>0$, we can find an $\eta>0$ such that tâtonnement converges linearly to an $\epsilon$-approximate equilibrium. Our results are achieved by leveraging the connection between tâtonnement and subgradient methods, and from there showing that the dual of the EG convex program satisfies the quadratic growth condition around its optimum. Moreover, we must then show bounded subgradients, which we establish by showing that the tâtonnement process is guaranteed to be bounded above and away from zero for its entire trajectory, assuming a sufficiently small stepsize. As is explained in Section 2, many existing works on tâtonnement use some modified version of the process where prices are generally kept away from zero either through projection to a region bounded away from zero, or by artificially upper-bounding excess demands. Our results show that for LFMs, such modifications are not necessary as long as stepsizes are sufficiently small. Finally, we extend our analysis framework to an important variant of LFMs: quasi-linear Fisher markets. For such markets, we again show that tâtonnement achieves a linear convergence rate to approximate equilibrium.
@@ 2 Related Work
Initially, economists studied the convergence of tâtonnement in continuous time. There, Arrow, Block, and Hurwicz (1959) proved that a continuous-time version of tâtonnement converges to ME for markets satisfying weak gross substitutability (WGS). In the remainder of the literature review, we focus on the discrete-time version of tâtonnement, which is the focus of our paper. Codenotti, McCune, and Varadarajan (2005) considered a discrete version of tâtonnement and proved that it converges to an approximate equilibrium in polynomial time for any exchange economy satisfying WGS. Though their tâtonnement updates are additive and similar to ours, they need to apply tâtonnement on a transformed market and use price projection in order to restrict prices to a region that is bounded and bounded away from zero. Secondly, they assume that demand is unique, which is not true for LFMs. Cole and Fleischer (2008) considered a multiplicative tâtonnement with artificially upper-bounded excess demands, and showed polynomial convergence for it on non-linear CES markets satisfying WGS ($10\leq\rho<$ 1). Cheung, Cole, and Rastogi (2012) extended this analysis to some non-WGS markets.
Cheung, Cole, and Devanur (2019) proved that the entropic tâtonnement is equivalent to generalized gradient descent with KL divergence, and thus is guaranteed to reach $\epsilon$ fraction of initial distance to ME measured in a potential function in $\mathcal{O}(\frac{1}{\epsilon})$ iterations for the Leontief Fisher market, and $\mathcal{O}(\log{\frac{1}{\epsilon}})$ iterations for the complementary CES Fisher market ($-\infty<\rho<0$). Avigdor-Elgrabli, Rabani, and Yadgar (2014); Cheung, Cole, and Rastogi (2012) considered multiplicative tâtonnement and showed convergence to an approximate equilibrium for Fisher markets with nested non-linear CES-type utilities.
Tâtonnement Papers Market Type Result
Additive tâtonnement $p^{t+1}_{j}=\Pi_{\Delta}\left(p^{t}_{j}+\eta z^{t}_{j}\right)$ Codenotti, McCune, and Varadarajan (2005) Exchange market WGS, unique demand Poly-time to approx ME
Multiplicative tâtonnement $p^{t+1}_{j}=p^{t}_{j}(1+\eta z^{t}_{j})$ Avigdor-Elgrabli, Rabani, and Yadgar (2014) Nested CES-Leontief $\rho\in(-\infty,0)\cup(0,1)$ $\mathcal{O}(\frac{1}{\epsilon^{3}})$
Multiplicative tâtonnement $p^{t+1}_{j}=p^{t}_{j}(1+\eta\min\{z^{t}_{j},1\})$ Cole and Fleischer (2008) Exchange market WGS $\mathcal{O}(E\log{\frac{1}{\epsilon}})$
Cheung, Cole, and Rastogi (2012) CES ($-1<\rho\leq 0$) Nested non-lin. CES $\mathcal{O}(\log{\frac{1}{\epsilon}})$
Entropic tâtonnement $p^{t+1}_{j}=p^{t}_{j}e^{\eta\min\{z^{t}_{j},1\}}$ Cheung, Cole, and Devanur (2019) CES ($-\infty\leq\rho<0$) $\mathcal{O}(\log{\frac{1}{\epsilon}})$ $\mathcal{O}(\frac{1}{\epsilon})\,\textnormal{if}\,\rho=-\infty$
Goktas, Zhao, and Greenwald (2023) CCNH $\mathcal{O}(\frac{1+E^{2}}{\epsilon})$
Entropic tâtonnement $p^{t+1}_{j}=\Pi_{\Delta}\big{(}p^{t}_{j}e^{\eta\min\{z^{t}_{j},1\}}\big{)}$ Cole and Tao (2019) Linear Large market assum. Linear conv. to approx ME
Additive tâtonnement $p^{t+1}_{j}=p^{t}_{j}+\eta z^{t}_{j}$ This work Linear Linear conv. to approx ME
Table 1: Comparison of results of different tâtonnement methods. The superscript denotes a modified version of the update rule. Note that only the last two papers consider linear Fisher markets. Cole and Fleischer (2008) and Goktas, Zhao, and Greenwald (2023) consider a broader class of utilities, but their parameter $E$ is infinity for linear utilities. $\Pi_{\Delta}(\cdot)$ denotes projection onto a bounded region. $E$ denotes the upper bound of the elasticity of demand.
@@
Goktas, Viqueira, and Greenwald (2021) studied the dual of an Eisenberg-Gale-like convex program from the perspective of the expenditure minimization problem (EMP). Goktas, Zhao, and Greenwald (2023) showed an $\mathcal{O}((1+E^{2})/T)$ convergence for an entropic tâtonnement process on Fisher markets with concave, continuous, nonnegative, and $1$-homogeneous  [^1] A utility function $u:\mathbb{R}^{m}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is $\alpha$-homogeneous if $u(ax)=a^{\alpha}u(x)$ for any $a\in\mathbb{R}_{\geq 0}$ and $x\in\mathbb{R}^{m}_{\geq 0}$. (CCNH) utilities, where $E$ denotes the upper bound of the elasticity of demand. Note that in the case of LFM, the elasticity of demand is unbounded, and thus this result gives an infinite upper bound.
@@
We summarize the above related work on the convergence of tâtonnement in Table 1. The table provides an overview of the different tâtonnement methods considered in the literature, along with a high-level description of convergence rates.
@@
Subgradient descent methods (SubGDs) have been extensively studied since their introduction in the 1960s. Historically, SubGD was only known to have an $O(1/\sqrt{T})$ convergence rate for the averaged iterate, where $T$ is the number of iterations. Moreover, this convergence was in terms of the optimality gap, and not in terms of the distance to the optimal solution. For tâtonnement, we are interested in the latter, as we want to know how quickly the prices converge to the market equilibrium prices. Recently, Zamani and Glineur (2023) showed a nearly $O(1/\sqrt{T})$ optimal rate for the last iterate, but again only in terms of optimality gap. It is known, however, that SubGD can achieve a faster convergence rate when the objective satisfies various notions of error-bound conditions (along with a bounded gradient condition). For prior work on SubGDs under error-bound conditions, see Johnstone and Moulin (2020) and references therein. Particularly relevant to us, SubGD is known to have linear convergence under the quadratic growth condition (QG) (Nedić and Bertsekas 2001; Karimi, Nutini, and Schmidt 2016), in terms of the distance to the optimal solution. This condition also allows one to connect convergence in optimality gap to convergence in iterates. We will show that this condition holds for LFMs, and thus SubGD has linear convergence in our setting.
@@ 3 Linear Fisher markets and Tâtonnement
We consider a linear Fisher market with $n$ buyers and $m$ divisible items. Each buyer $i$ has a budget $B_{i}>0$ and valuation $v_{ij}\geq 0$ for each item $j$. Denote $B=\sum_{i\in[n]}B_{i}$ and $B_{\min}=\min_{i\in[n]}B_{i}$. We denote $v_{i}=(v_{i1},\dots,v_{im})$ for all $i$ and $v$ as the valuation matrix whose $i$-th row is $v_{i}$. An allocation (or bundle) of items $x_{i}\in\mathbb{R}_{\geq 0}^{m}$ gives buyer $i$ a utility of $u_{i}(x_{i})=\langle v_{i},x_{i}\rangle$. Without loss of generality, we assume the following:
@@
$\mathcal{D}_{i}(p):=\operatorname\textup{*}{\textnormal{Argmax}}_{x_{i}}\left\{\langle v_{i},x_{i}\rangle:\langle p,x_{i}\rangle\leq B_{i},\,x_{i}\geq 0\right\}.$ (3)
@@
A Fisher market is said to reach a market equilibrium if allocations $x^{\textup{*}}_{i}\in\mathbb{R}^{m}_{\geq 0},\;\forall\;i\in[n]$ and prices $p^{\textup{*}}\in\mathbb{R}^{m}_{\geq 0}$ satisfy
  • Buyer optimality: $*x^{\textup{*}}_{i}\in\mathcal{D}_{i}(p^{\textup{*}})$ for all $i\in[n]$,
  • Market clearance: $\sum_{i=1}^{n}x^{\textup{*}}_{ij}=1$ for all $j\in[m]$.
@@
$\begin{array}[]{rll}\bullet&\text{Pick}\;x_{i}^{t}\in\mathcal{D}_{i}(p^{t})&\forall\;i\in[n]\\[3.0pt] \bullet&z^{t}_{j}=\sum_{i=1}^{n}x_{i}^{t}-1&\forall\;j\in[m]\\[3.0pt] \bullet&p^{t+1}_{j}=p^{t}_{j}+\eta^{t}z^{t}_{j}&\forall\;j\in[m].\end{array}$ (4)
Typically, many authors have considered some form of projection to ensure $p^{t+1}\geq 0$, or even such that it is strictly bounded away from zero. We will show that for LFMs, it is possible to select $\eta^{t}$ such that the prices are ensured to be strictly bounded away from zero without projection. This is critical for our later convergence results. We will focus on the case where stepsizes are constant, i.e. $\eta^{t}=\eta$ for some $\eta>0$.
As shown in Section 1, Eq. 4 is equivalent to applying subgradient descent method on the following non-smooth convex minimization problem:
@@
In particular, starting with some feasible point $p^{0}\in\mathbb{R}^{m}_{\geq 0}$, at time $t$, we update $p$ as follows:
$\begin{array}[]{rl}\bullet&\text{Pick}\;g(p^{t})\in\partial\varphi(p^{t})\\[3.0pt] \bullet&p^{t+1}=\Pi_{\mathbb{R}^{m}_{\geq 0}}\left(p^{t}-\eta g(p^{t})\right).\end{array}$ (6)